Peter sold pumpkins from his farm. He sold jumbo pumpkins for $9.00 each, and he sold regular pumpkins for $4.00 each. peter sold 80 pumpkins and collected $395.00.
How many jumbo pumpkins and regular pumpkins did he sell?
Please tell me what the word "each" stands for in word problems.

Answer by rapaljer(4551)

The word "each" means the price of each pumpkin in that particular category.

The equation is the sum of the cost of EACH type of pumpkin times the number of pumpkins sold at EACH price.

Let x = number of jumbo pumpkins at $9 EACH. The rest of the 80 pumpkins were regular $4 pumpkins, so this means that

80-x = number of regular pumpkins at $4 EACH.

The equation is
9*x + 4*(80-x) = 395
9x + 320 - 4x = 395
5x + 320=395
5x= 75
x= 15 Jumbo Pumpkins
80-x = number of regular pumpkins
80-15= 65 Regular Pumpkins

Check: See if this adds up to $395
15*$9 + 65*$4
$135 + $260= $395 It checks!!

By the way, if you need help with ALL kinds of Word Problems, especially this type of problem, please see my own website. Do a "Bing" search for my last name "Rapalje", and look for "Rapalje Homepage" which should be at the top of the list. Near the top of my Homepage, look for the link "Basic Intermediate and College Algebra: One Step at a Time." Choose "Basic Algebra", and look in Chapter 1 for the topics of Word Problems. This is a free user-friendly explanation, with examples and exercises. See also the "MATH IN LIVING COLOR" pages that go with these pages, where many of these exercises are solved in COLOR!

R^2